# Hydrodynamic memory and Quincke rotation

### v1

@@ VII Conclusion
A mathematical model for the Quincke rotation of a spherical particle that accounts for the inertia of the particle and the unsteady Stokes flow in the surrounding fluid is presented and analyzed. This model takes the form of an integro-differential, infinite-dimensional dynamical system that we term the ‘memory Lorenz equations’ (MLE). The primary conclusion is that the linearized fluid inertia, or hydrodynamic memory as we call it, causes an increase in the threshold field strength for observation of chaotic rotor dynamics. This is in qualitative agreement with previous experimental observations. A secondary conclusion is that increasing the ratio of momentum diffusion to dipole relaxation times ($\gamma$) leads to an increasing range of field strengths for which multi-stability between steady and chaotic rotation exists. A simplified form of the MLE (i.e., sMLE), which may be relevant to rotor dynamics in non-Newtonian fluids, indicates that the above conclusions could be a generic feature of Lorenz-type systems with distributed memory. This method of suppressing the onset of chaos with memory effects may be relevant to the field of chaos control [33, 34] where time-delayed feedback is routinely used to control chaotic dynamics in various systems. Further, since the effects of exponentially fading memory for sMLE are captured by just one additional ODE coupled to the original $3$D Lorenz equations, this dynamical system might provide useful insights into suitable chaos control methods using simple feedback mechanisms [35].
@@
It would, of course, be desirable to test in more detail these predictions against fresh experiments. Since the MLE model appears to preserve all the intricate dynamical features of LE, experiments with a Quincke rotor might form a testing bed to experimentally realize the complex bifurcations associated with Lorenz chaos. Natural extensions of the present work include rotor dynamics in time-dependent fields, or dynamics in the presence of a background shear flow.

### v2

@@ VIII Conclusion
A mathematical model for the Quincke rotation of a spherical particle that accounts for the inertia of the particle and the unsteady Stokes flow in the surrounding fluid is presented and analyzed. This model takes the form of an integro-differential, infinite-dimensional dynamical system that we term the ‘memory Lorenz equations’ (MLE). The primary conclusion is that the linearized fluid inertia, or hydrodynamic memory as we call it, causes an increase in the threshold field strength for observation of chaotic rotor dynamics. This is in qualitative agreement with previous experimental observations. A secondary conclusion is that increasing the ratio of momentum diffusion to dipole relaxation times ($\gamma$) leads to an increasing range of field strengths for which multi-stability, or co-existence, between steady and chaotic rotation occurs. In the experiments of [20], the radius of the cylindrical rotor (a glass tube) is around 1mm, whereas the spacing between the electrodes that supply the electric field is around 1cm. The rotor is placed at the midpoint of the electrodes. Hence, a more accurate theoretical treatment would require accounting for the hydrodynamic interactions between the rotor and the electrodes, which would alter the form of the memory kernel in the MLE. However, we do not expect the primary and secondary conclusions to be affected by our neglect of such interactions. Further, spherical particles often undergo Quincke rotation near a wall (upon which they have settled), in which case they are referred to as ‘Quincke rollers,’ since the rotation is accompanied by translational motion along the wall. Certainly, the hydrodynamic interaction between the particle and wall would affect the hydrodynamic memory in this case.
A simplified form of the MLE (i.e., sMLE), which may be relevant to rotor dynamics in non-Newtonian fluids, indicates that the above conclusions could be a generic feature of Lorenz-type systems with distributed memory. This method of suppressing the onset of chaos with memory effects may be relevant to the field of chaos control [33, 34] where time-delayed feedback is routinely used to control chaotic dynamics in various systems. Further, since the effects of exponentially fading memory for sMLE are captured by just one additional ODE coupled to the original 33D Lorenz equations, this dynamical system might provide useful insights into suitable chaos control methods using simple feedback mechanisms [35].
@@
It would, of course, be desirable to test in more detail our predictions against fresh experiments; for example, can the co-existence between steady and chaotic rotation be observed experimentally? Since the MLE model appears to preserve all the intricate dynamical features of LE, experiments with a Quincke rotor might form a testing bed to experimentally realize the complex bifurcations associated with Lorenz chaos. Natural extensions of the present work include rotor dynamics in time-dependent fields, or dynamics in the presence of a background shear flow.